An investor invested a total of 2400.oo in two mutual funds. One fund earned a 4%and the other earned a 2% profit. If the investor's total profit was 86.00, how much was invested in each mutual fund?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y
2400
Interest----- 86

Part I 4.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + 1 y= 2400 -------------1
Interest on both investments
4.00% x + 2.00% y= 86
Multiply by 100
4 x + 2 y= 8600.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -9600.00
Add this to (2)
0 x -2 y= -1000
divide by -2
y = 500
Part I 4.00% $ 1900
Part II 2.00% $ 500

CHECK
1900 --------- 4.00% ------- 76.00
500 ------------- 2.00% ------- 10.00
Total -------------------- 86.00

m.ananth@hotmail.ca